# Counting collisions in random circuit sampling for benchmarking quantum computers

### v1

@@ Abstract
We show that counting the number of collisions (re-sampled bitstrings) when measuring a random quantum circuit provides a practical benchmark for the quality of a quantum computer and a quantitative noise characterization method. We analytically estimate the difference in the expected number of collisions when sampling bitstrings from a pure random state and when sampling from the classical uniform distribution. We show that this quantity, if properly normalized, can be used as a collision anomaly benchmark or as a collision volume test which is similar to the well-known quantum volume test, with advantages (no classical computing cost) and disadvantages (high sampling cost). We also propose to count the number of cross-collisions between two independent quantum computers running the same random circuit in order to obtain a cross-validation test of the two devices. Finally, we quantify the sampling cost of quantum collision experiments. We find that the sampling cost for running a collision volume test on state-of-the-art processors (e.g. 20 effective clean qubits) is quite small: less than $10^{5}$ shots. For large-scale experiments in the quantum supremacy regime the required number of shots for observing a quantum signal in the observed number of collisions is currently infeasible ($>10^{12}$), but not completely out of reach for near-future technology.
@@ I Introduction
In this work we propose a remarkably simple validation protocol that can be summarized in a single sentence: repeatedly execute a random quantum circuit and count the number of collisions, i.e., the number of times a bitstring, that was previously sampled, is sampled again.111 There can be different ways of defining the number of collisions. Given $k$ equal samples, one can count all the $k(k-1)/2$ equal pairs, or the number of $k-1$ re-sampling events . In this work we use the latter approach. In any case, both definitions become equivalent in the rare collision limit $k\leq 2$. It turns out that the expected number of collisions is larger for a good quantum computer compared to a noisy one, or compared to the extreme limit of a (classical) random number generator. By comparing the observed number of collisions with the theoretical ideal value, one can quantitatively deduce the fidelity of the quantum computation. More practically, the same experimental procedure can be used to obtain a collision volume (CV) benchmark which, like the standard quantum volume metric cross2019validating ; blume2020volumetric ; baldwin2022re , quantifies the maximum number of clean effective qubits bishop2017quantum ; moll2018quantum ; kechedzhi2023effective ; larose2022error available in the quantum device.
Figure 1: Numerical simulation of the number of collisions $\hat{R}$ with respect to the number of samples $N$, for a $16$-qubit random pure quantum state (blue) and for the uniform distribution over bitstrings of $n$=16 bits (orange). The theoretical lines correspond to Eqs. (2) and (II). The vertical line is the educated guess for the number of shots required to observe $\gtrsim 500$ collisions, as derived in Eq. (6). $D$ is the dimension of the sample space which, in this example, is $2^{16}$.
@@
Unfortunately there is no free lunch since the exponential classical simulation cost required by the QV and XEB methods is pushed, by the quantum collision protocol, to the quantum side. Indeed, in order to observe a nonzero number of collision events when sampling a random state of $n$ qubits, one needs a number of shots larger than $2^{n/2}$. This is an exponential cost, even though, it is still much smaller than the $\mathcal{O}(2^{n})$ cost of a brute-force classical simulation. Despite the exponential asymptotic scaling of the sampling cost, the potential advantage of the method proposed in this work is given by its operational simplicity and, perhaps, by the practical convenience of just taking more shots instead of designing and running exponentially hard classical simulations. It is worth noting that standard validation methods for large scale computations (close to the quantum supremacy regime) require advanced HPC classical simulations which are technically involved, economically expensive, and energetically inefficient. Therefore, the possibility of trading classical simulation cost for quantum sampling cost opens up a new interesting perspective.
@@
We also propose to use a similar collision-based protocol for the cross-validation of two independent quantum computers: by counting the number of cross-collisions observed when sampling two quantum computers running the same random circuit, one can estimate the joint quality of both devices. Similarly, one can use a reliable and trusted quantum computer to validate an untrustworthy device. The number of collisions observed for single device provides a good noise characterization benchmark only if used in a fair and careful experiment. In fact, the benchmark could be intentionally or unintentionally spoofed by any mock device that samples from a sufficiently non-uniform distribution, even if unrelated to the correct quantum state. Quite interestingly instead, the cross-validation test between two independent quantum computers is much more difficult to spoof, as discussed in Sections IV and VII.
@@ II Expected number of collisions
where, in the last step, we used $(1-1/D)^{D}\simeq e^{-1}$ which is a good approximation for large $D$. For $N/D\ll 1$, the expected collision frequency can be expanded as follows222 Neglecting higher order terms in $N/D$ is not always a good approximation. We recommend avoiding the first-order approximation when looking for an accurate estimate. :
@@
[…]
@@
(iii) The number of collisions can be used to distinguish the measurement distribution of a random pure state (e.g. the results of an ideal quantum computer) from the uniform distribution (e.g. the results of a very noisy quantum computer or of a classical pseudo-random number generator). In the limit $N/D\rightarrow 0$ the number of quantum collisions is twice the number of classical collisions. This fact is consistent with the similar doubling behavior of the linear XEB and of the collision probability,333 In general, the expected number of collisions is not directly proportional to the collision probability. From Eq. (1), it is evident that the former depends on all the moments $\sum_{j}p_{j}^{k}$ of the probability distribution, while the latter corresponds to $\sum_{j}p_{j}^{2}$. see e.g. Ref. aaronson2019classical and Theorems 2.10 and 3.5 of Ref. liu2021moments . The gap is smaller for finite values of $N/D$, e.g., for $D=2^{16}$ and $N=32\sqrt{D}\simeq 8.2\times 10^{3}$, we get $\mathbb{E}_{q}(\hat{R})/\mathbb{E}_{u}(\hat{R})\simeq 1.85$.
@@ III A collision volume test
which quantifies the deviation of the observed number of collisions $\hat{R}$ from the base value associated to the uniform distribution. By construction, the expected value of $\hat{\Delta}$ is $1$ when sampling from a typical random quantum state, while it is 00 when sampling from the uniform distribution.
@@ IV Cross-validation of two quantum computers
In this section we are interested in the expected number of cross-collisions when sampling from two different probability distributions corresponding, e.g., to two different quantum computers operated by independent agents: Alice and Bob. Let $\{p_{j}^{(A)}:j=1\dots D\}$ and $\{p_{j}^{(B)}:j=1\dots D\}$ be the probability distributions of Alice and Bob, respectively. If Alice samples $N_{A}$ bitstrings, and Bob samples $N_{B}$ bitstrings, what is the expected number of cross-collisions? There can be slightly different ways of defining the number of cross collisions. Here we choose a definition that allows us to recycle most of the previous theory developed for a single device.
@@
Figure 3: Numerical simulation of the number of cross-collisions $\hat{R}_{X}$ with respect to the number of samples $N$. We consider the cases in which Alice and Bob sample, with the same number of shots, from the same random quantum states (blue), from the uniform distribution (orange), from a random quantum state and the uniform distribution (green). The theoretical curves correspond to Eqs. (10), (11) and (IV).
@@
Figure 4: Numerical simulation of the cross-collision anomaly $\hat{\Delta}_{X}$ defined in Eq. (7). We consider the cases in which Alice and Bob sample with the same number of shots from: the same random quantum states (blue), from the uniform distribution (orange), from a random quantum state and the uniform distribution (green), from different random states (red).
@@
So, in the limit $D\rightarrow\infty$, when both Alice and Bob sample from the same quantum state the number of cross-collisions is twice the value of the other cases involving the uniform distribution. We conjecture that $\mathbb{E}_{qq^{\prime}}(\hat{R}_{X})\leq\mathbb{E}_{uu}(\hat{R}_{X})<\mathbb{E}_{qq}(\hat{R}_{X})$, where $q$ and $q^{\prime}$ are associated to different pure random states. In other words, Alice and Bob get more cross-collisions if they both sample from the same quantum sate with sufficiently good fidelity. In this sense, the observation of a large number of collisions is a good validation test for the joint quality of both devices. Moreover, if Alice has a sufficiently good and trusted quantum computer, she can use it as a reference for testing the quality of an untrusted quantum computer (Bob’s device), by simply measuring the number of cross-collisions without performing any classical simulation.
@@ V Expected number of collisions for a quantum device subject to depolarizing noise
In the previous sections we studied the expected number of collisions involving pure states, i.e., the results of a noiseless quantum computer. Unfortunately all existing quantum computers are noisy. What happens to the expected number of collisions if a quantum computer is noisy?
It has been numerically and experimentally demonstrated boixo2018characterizing ; arute2019quantum , that the quantum state generated by a noisy quantum computer after the execution of a random circuit is well approximated by the following density matrix:
@@
Figure 6: Numerical simulation of the quantum collision anomaly $\hat{\Delta}$ defined in Eq. (7), for a $16$-qubit random quantum state subject to depolarizing noise with fidelity parameter $\alpha$ as defined in Eq. (15).
Assume that we have two fictitious agents, Alice who measures with a perfect quantum computer $N_{A}=\alpha N$ bitstrings from $|\psi\rangle$ and Bob who measures $N_{B}=N(1-\alpha)$ bitstrings from the uniform distribution. In the limit of large N, if they merge and randomly shuffle their results, they effectively obtain $N$ samples from the noisy quantum state in Eq. (15). Moreover, if we are just interested in counting the number of collisions, the shuffling operation is irrelevant and we can consider the bitstrings of Alice and Bob as distinct datasets. But this is exactly the situation that we have studied in the previous section, where we derived Eq. (9). More explicitly, if $q_{\alpha}$ is the probability distribution associated to the noisy quantum state (15), the expected number of collisions is:
@@
corresponding to a constant floor of $0.5N/D$ (that of the uniform distribution) plus an additive quantum signal scaling as $\alpha^{2}$ which is the deviation that we are interested in measuring. Assuming that statistical fluctuations depend weakly on $\alpha$, we can counteract the effect of noise by scaling $N$ as $N\rightarrow\alpha^{-1}N$ in order to obtain a signal comparable to the noiseless case. Accordingly, we update the empirical educated guess given in Eq. (6) with a new formula which takes noise into account:
@@ VI Sampling cost: from the collision volume regime up to the quantum supremacy regime
In Fig. 7 we give a cost estimate in terms of the number of shots and in terms of the execution time in both regimes: the collision volume regime (few qubits) and the quantum supremacy regime (many qubits). We observe that, running a collision volume test for state-of-the-art NISQ computers ($n\approx 20$ effective qubits with $\alpha>0.5$) is quite easy since $N\leq 10^{5}$ shots are enough. On the other hand, for validating quantum computations in the quantum supremacy regime ($n>53$ and $\alpha<0.002$) by empirically estimating $\hat{\Delta}\simeq\alpha^{2}>0$ or $\hat{\Delta}_{X}\simeq\alpha^{2}>0$ with sufficient statistical confidence, at least $N>10^{12}$ shots are required. Assuming a measurement repetition rate of 1 MHz, this corresponds to more than 3 weeks of sampling time (see vertical line in Fig. 7). Running a weeks-long quantum experiment is arguably infeasible with current technology, but it may become feasible with future technology.
@@
For example, with $M$ independent processors (e.g. in a multi-core QPU), one can reduce the total time by a factor of $1/M$. Moreover, it is reasonable to expect fidelity improvements, from $\alpha\approx 0.002$ as estimated in Ref. arute2019quantum to larger values. Depending on the magnitude of such technological improvements, the large quantum sampling cost required to validate quantum computations by measuring quantum collisions could become competitive with respect to the intensive classical simulation cost required by classical validation strategies boixo2018characterizing ; arute2019quantum . The potential advantage of the collision-based validation method is perhaps even stronger if instead of the time cost we consider the energy consumption. Indeed, the experimental task quantum circuit sampling, even if technically demanding, can be more energy efficient than running large classical simulations on supercomputers scholes2010green ; villalonga2020establishing ; auffeves2022quantum .

### v2

@@ Abstract
We show that counting the number of collisions (re-sampled bitstrings) when measuring a random quantum circuit provides a practical benchmark for the quality of a quantum computer and a quantitative noise characterization method. We analytically estimate the difference in the expected number of collisions when sampling bitstrings from a pure random state and when sampling from the classical uniform distribution. We show that this quantity, if properly normalized, can be used as a collision anomaly benchmark or as a collision volume test which is similar to the well-known quantum volume test, with advantages (no classical computing cost) and disadvantages (high sampling cost). We also propose to count the number of cross-collisions between two independent quantum computers running the same random circuit in order to obtain a cross-validation test of the two devices. Finally, we quantify the sampling cost of quantum collision experiments. We find that the sampling cost for running a collision volume test on state-of-the-art processors (e.g. 20 effective clean qubits) is quite small: less than $10^{5}$ shots. For large-scale experiments in the quantum supremacy regime, the required number of shots for observing a quantum signal in the observed number of collisions is currently infeasible ($>10^{12}$), but not completely out of reach for near-future technology.
@@ I Introduction
In this work we propose a remarkably simple validation protocol that can be summarized in a single sentence: repeatedly execute a random quantum circuit and count the number of collisions, i.e., the number of times a bitstring, that was previously sampled, is sampled again footnote1 . It turns out that the expected number of collisions is larger for a good quantum computer compared to a noisy one, or compared to the extreme limit of a (classical) random number generator. By comparing the observed number of collisions with the theoretical ideal value, one can quantitatively deduce the fidelity of the quantum computation. More practically, the same experimental procedure can be used to obtain a collision volume (CV) benchmark which, like the standard quantum volume metric cross2019validating ; blume2020volumetric ; baldwin2022re , quantifies the maximum number of clean effective qubits bishop2017quantum ; moll2018quantum ; kechedzhi2023effective ; larose2022error available in the quantum device.
Figure 1: Numerical simulation of the number of collisions $\hat{R}$ with respect to the number of samples $N$, for a single $16$-qubit random pure quantum state (blue) and for the uniform distribution over bitstrings of $n$=16 bits (orange). The theoretical lines correspond to Eqs. (2) and (II). The vertical line is the educated guess for the number of shots required to observe $\gtrsim 500$ collisions, as derived in Eq. (6). $D$ is the dimension of the sample space which, in this example, is $2^{16}$.
@@
Benchmarks based on random circuits Classical cost Quantum cost Incoherent errors Coherent errors Classical spoofing hardness
QV blume2020volumetric and linear XEB boixo2018characterizing ; aaronson2019classical exponential polynomial sensible sensible Hard aaronson2019classical with ongoing debate barak2020spoofing
Collision test (this work) none exponential sensible to most insensible Easy
Cross-collision test (this work) none exponential sensible sensible Unknown (assuming no cooperation)
Table 1: Qualitative summary of the main properties of collision-based benchmarks and comparison with other existing methods. By sensible to most errors we mean that the collision test executed on a single device is sensible to most incoherent errors in a real-world scenario, but one can easily find ad-hoc counterexamples (e.g. a noise model that, with some probability, replaces the output state of the circuit with a fixed product state). We remark that if Alice and Bob can agree on a common cheating strategy (e.g. on a secret set of shared bitstrings) the cross-collision test can be trivially spoofed. On the contrary, assuming Alice and Bob cannot adversarially cooperate, we conjecture that the test is as hard to spoof as linear XEB.
@@
Unfortunately, there is no free lunch since the exponential classical simulation cost required by the QV and XEB methods is pushed, by the quantum collision protocol, to the quantum side. Indeed, in order to observe a nonzero number of collision events when sampling a random state of $n$ qubits, one needs a number of shots larger than $2^{n/2}$. This is an exponential cost, even though, it is still much smaller than the $\mathcal{O}(2^{n})$ cost of a brute-force classical simulation. Despite the exponential asymptotic scaling of the sampling cost, the potential advantage of the method proposed in this work is given by its operational simplicity and, perhaps, by the practical convenience of just taking more shots instead of designing and running exponentially hard classical simulations. It is worth noting that standard validation methods for large-scale computations (close to the quantum supremacy regime) require advanced HPC classical simulations which are technically involved, economically expensive, and energetically inefficient. Therefore, the possibility of trading classical simulation cost for quantum sampling cost opens up a new and interesting perspective.
An important aspect to consider is that the number of collisions observed for a single device provides a good noise characterization benchmark only if most errors are incoherent and if the experiment is fair (non-advarsarial). Indeed, the benchmark could be intentionally or unintentionally spoofed by any mock device that samples from a sufficiently non-uniform distribution, even if unrelated to the correct quantum state. For the same reason, systematic coherent errors such as miscalibration of gates are undetectable by counting collisions on a single device, since a wrong pure random state gives the same collision statistics as the correct pure random state. This implies that the number of observed collisions in a single quantum computer should not be used as a conclusive and absolute validation benchmark, but as a simple and practical diagnostic test augmenting the toolkit of existing benchmarks. In Table 1 we summarize the main properties of collision-based benchmarks and we compare them with other randomized methods.
@@
We also propose to use a similar collision-based protocol for the cross-validation of two independent quantum computers: by counting the number of cross-collisions observed when sampling two quantum computers running the same random circuit, one can estimate the joint quality of both devices. Similarly, one can use a reliable and trusted quantum computer to validate an untrustworthy device. Quite interestingly, compared to the single-device case, the cross-validation test between two independent quantum computers is much more difficult to spoof, as discussed in Sections IV and VII. Moreover, the cross-collision benchmark is also sensitive to coherent errors under the reasonable assumption that the noise models of the two quantum computers are statistically independent.
@@ II Expected number of collisions
where, in the last step, we used $(1-1/D)^{D}\simeq e^{-1}$ which is a good approximation for large $D$. For $N/D\ll 1$, the expected collision frequency can be expanded as follows footnote2 :
@@
where the approximation in the second line is valid for $D\gg 1$ since, in this regime, the term $e^{-Dp}$ exponentially suppresses the integrand for large values of $p$ such that the approximation $(1-p)^{1/p}\simeq e^{-1}$ is justified and the upper integration limit can be extended to +$\infty$+ ∞. For $N/D\ll 1$, the expected collision frequency can be expanded as follows footnote2 :
@@
(iii) The number of collisions can be used to distinguish the measurement distribution of a random pure state (e.g. the results of an ideal quantum computer) from the uniform distribution (e.g. the results of a very noisy quantum computer or of a classical pseudo-random number generator). In the limit $N/D\rightarrow 0$ the number of quantum collisions is twice the number of classical collisions. This fact is consistent with the similar doubling behavior of the linear XEB and of the collision probability footnote3 , see e.g. Ref. aaronson2019classical and Theorems 2.10 and 3.5 of Ref. liu2021moments . The gap is smaller for finite values of $N/D$, e.g., for $D=2^{16}$ and $N=32\sqrt{D}\simeq 8.2\times 10^{3}$, we get $\mathbb{E}_{q}(\hat{R})/\mathbb{E}_{u}(\hat{R})\simeq 1.85$.
@@ III A collision volume test
which quantifies the deviation of the observed number of collisions $\hat{R}$ from the base value associated to the uniform distribution. By construction, the expected value of $\hat{\Delta}$ is $1$ when sampling from a typical random quantum state, while it is 00 when sampling from the uniform distribution. See Fig. 2 for a numerical simulation.
@@ IV Cross-validation of two quantum computers
In this section, we are interested in the expected number of cross-collisions when sampling from two different probability distributions corresponding, e.g., to two different quantum computers operated by independent agents: Alice and Bob. Let $\{p_{j}^{(A)}:j=1\dots D\}$ and $\{p_{j}^{(B)}:j=1\dots D\}$ be the probability distributions of Alice and Bob, respectively. If Alice samples $N_{A}$ bitstrings, and Bob samples $N_{B}$ bitstrings, what is the expected number of cross-collisions? There can be slightly different ways of defining the number of cross-collisions. Here we choose a definition that allows us to recycle most of the previous theory developed for a single device.
@@
Figure 3: Numerical simulation of the number of cross-collisions $\hat{R}_{X}$ with respect to the number of samples $N$. We consider the cases in which Alice and Bob sample, with the same number of shots, from the same (16-qubit) random quantum states (blue), from the uniform distribution (orange), from a random quantum state and the uniform distribution (green). The theoretical curves correspond to Eqs. (10), (11) and (IV).
Figure 4: Numerical simulation of the cross-collision anomaly $\hat{\Delta}_{X}$ defined in Eq. (7). We consider the cases in which Alice and Bob sample with the same number of shots from: the same 16-qubit random quantum states (blue), from the uniform distribution (orange), from a random quantum state and the uniform distribution (green), and from different random states (red).
@@
So, in the limit $D\rightarrow\infty$, when both Alice and Bob sample from the same quantum state the number of cross-collisions is twice the value of the other cases involving the uniform distribution. From numerical evidence (see e.g. Figs. 3 and 4) and from intuitive arguments footnote4 , we conjecture that $\mathbb{E}_{qq^{\prime}}(\hat{R}_{X})\leq\mathbb{E}_{uu}(\hat{R}_{X})<\mathbb{E}_{qq}(\hat{R}_{X})$, where $q$ and $q^{\prime}$ are associated to different pure random states. In other words, Alice and Bob get more cross-collisions if they both sample from the same quantum state with sufficiently good fidelity. In this sense, the observation of a large number of collisions is a good validation test for the joint quality of both devices. Moreover, if Alice has a sufficiently good and trusted quantum computer, she can use it as a reference for testing the quality of an untrusted quantum computer (Bob’s device), by simply measuring the number of cross-collisions without performing any classical simulation.
@@ V Expected number of collisions for a quantum device subject to depolarizing noise
In the previous sections, we studied the expected number of collisions involving pure states, i.e., the results of a noiseless quantum computer. Unfortunately, all existing quantum computers are noisy. What happens to the expected number of collisions if a quantum computer is noisy?
It has been numerically and experimentally demonstrated boixo2018characterizing ; arute2019quantum (see also dalzell2024random for theoretical arguments), that the quantum state generated by a noisy quantum computer after the execution of a random circuit is typically well approximated by the following density matrix:
@@
Figure 6: Numerical simulation of the quantum collision anomaly $\hat{\Delta}$ defined in Eq. (7), for a $16$-qubit random quantum state subject to global depolarizing noise with fidelity parameter $\alpha$ as defined in Eq. (15).
Assume that we have two fictitious agents, Alice who measures with a perfect quantum computer $N_{A}=\alpha N$ bitstrings from $|\psi\rangle$ and Bob who measures $N_{B}=N(1-\alpha)$ bitstrings from the uniform distribution. In the limit of large N, if they merge and randomly shuffle their results, they effectively obtain $N$ samples from the noisy quantum state in Eq. (15). Moreover, if we are just interested in counting the number of collisions, the shuffling operation is irrelevant and we can consider the bitstrings of Alice and Bob as distinct datasets. But this is exactly the situation that we have studied in the previous section, where we derived Eq. (9). More explicitly, if $q_{\alpha}$ is the probability distribution associated with the noisy quantum state (15), the expected number of collisions is:
@@
corresponding to a constant floor of $0.5N/D$ (that of the uniform distribution) plus an additive quantum signal scaling as $\alpha^{2}$ which is the deviation that we are interested in measuring. Assuming that statistical fluctuations depend weakly on $\alpha$, we can counteract the effect of noise by scaling $N$ as $N\rightarrow\alpha^{-1}N$ in order to obtain a signal comparable to the noiseless case. Accordingly, we update the empirical educated guess given in Eq. (6) with a new formula that takes noise into account:
@@ VI Sampling cost: from the collision volume regime up to the quantum supremacy regime
In Fig. 7 we give a cost estimate in terms of the number of shots and in terms of the execution time in both regimes: the collision volume regime (few qubits) and the quantum supremacy regime (many qubits). We observe that running a collision volume test for state-of-the-art NISQ computers ($n\approx 20$ effective qubits with $\alpha>0.5$) is quite easy since $N\leq 10^{5}$ shots are enough. On the other hand, for validating quantum computations in the quantum supremacy regime ($n>53$ and $\alpha<0.002$) by empirically estimating $\hat{\Delta}\simeq\alpha^{2}>0$ or $\hat{\Delta}_{X}\simeq\alpha^{2}>0$ with sufficient statistical confidence, at least $N>10^{12}$ shots are required. Assuming a measurement repetition rate of 1 MHz, this corresponds to more than 3 weeks of sampling time (see vertical line in Fig. 7). Running a weeks-long quantum experiment is arguably infeasible with current technology, but it may become feasible with future technology.
@@
For example, with $M$ independent processors (e.g. in a multi-core QPU), one can reduce the total time by a factor of $1/M$. Moreover, it is reasonable to expect fidelity improvements, from $\alpha\approx 0.002$ as estimated in Ref. arute2019quantum to larger values. Depending on the magnitude of such technological improvements, the large quantum sampling cost required to validate quantum computations by measuring quantum collisions could become competitive with respect to the intensive classical simulation cost required by classical validation strategies boixo2018characterizing ; arute2019quantum . The potential advantage of the collision-based validation method is perhaps even stronger if, instead of the time cost, we consider the energy consumption. Indeed, the experimental task quantum circuit sampling, even if technically demanding, can be more energy efficient than running large classical simulations on supercomputers scholes2010green ; villalonga2020establishing ; auffeves2022quantum .